OpenStudy (anonymous):

Find the slope and y-intercept for each equation and graph the equation: y= 3/2x

OpenStudy (anonymous):

\[y= \frac{ 3 }{ 2 } \]

ganeshie8 (ganeshie8):

y = mx + c c = y intercept m = slope

OpenStudy (anonymous):

so how would i do tht now tht i have the formula

ganeshie8 (ganeshie8):

compare both

ganeshie8 (ganeshie8):

y = mx + c y = \(\frac{3}{2}x + 0\) m = ? c = ?

OpenStudy (anonymous):

can u draw it please